OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2). −9 −7 −10 7

OpenStudy (anonymous):

f(x) = −4x − 1---just plug in f(-2) = −4(-2) − 1

OpenStudy (anonymous):

f(-2) = 7 f(-9) = 35 f(-7) = 27 f(-10) = 39 f(7) = -29

OpenStudy (anonymous):

how do you do that ?

OpenStudy (anonymous):

If you are doing f(-2), you just plug -2 into every x in the equation. Ex: f(-2) = -4(-2) - 1 f(-2) = 8 - 1 f(-2) = 7